Calculate the rate of heat conduction out of the human body, assuming that the core internal temperature is 37.31 ºC, the skin temperature is 33.49 ºC, the thickness of the tissues between averages 1.46 cm, and the surface area is 1.27 m2. (skin = 0.37W/m⋅°C)


1
Expert's answer
2022-05-09T14:49:52-0400
"P=\kappa\frac{\Delta T}{\Delta x}A"

"P=0.37*\frac{37.31-33.49}{0.0146}*1.27=123\:\rm W"